As purchasing agent for Eynan Enterprises in​ Richmond, Virginia, you ask your buyer to provide you with a ranking of​ "excellent" (worth 4​ points), "good"​ (3 points),​ "fair" (2​ points), or​ "poor" ​ (1 point) for a variety of characteristics for two potential​ vendors, Donna Inc.​ (D) and Kay Corp.​ (K). You suggest that​ "Products" total be weighted 0.4​,

the​ "Company" total be weighted 0.2​, and the​ "Service" and​ "Sales" totals each be weighted 0.2. The buyer has returned the following ranking. Which of the two vendors would you​ select?

Table​ 1: Company

Excellent​ (4)

Good​ (3)

Fair​ (2)

Poor​ (1)

Financial Strength

nothing

nothing

KD

Manufacturing Range

nothing

nothing

K

D

Research Facilities

K

D

Geographical Locations

K

nothing

D

nothing

Management

nothing

KD

nothing

nothing

Labor Relations

nothing

nothing

nothing

DK

Trade Relations

nothing

nothing

K

D

Table​ 2: Service

Excellent​ (4)

Good​ (3)

Fair​ (2)

Poor​ (1)

Deliveries on Time

KD

Handling of Problems

KD

Technical Assistance

K

D

Table​ 3: Products

Excellent​ (4)

Good​ (3)

Fair​ (2)

Poor​ (1)

Quality

K

nothing

D

nothing

Price

nothing

nothing

D

K

Packing

nothing

nothing

K

D

Table​ 4: Sales

Excellent​ (4)

Good​ (3)

Fair​ (2)

Poor​ (1)

Product knowledge

D

K

Sales Calls

nothing

nothing

nothing

DK

Sales Service

K

D

You should choose vendor with a weighted score of

nothing. ​(Enter your response rounded to one decimal​ place.)

please show steps and answer
Answer #1

Table for Ratings of both vendors are shown below to compute the totals under each head, namely, company, service, sales and product

Table 1: Company Kay Corp Donna Inc
Financial Strength 1 1
Manufacturing Range 2 1
Research Facilities 4 2
Geographical Locations 4 2
Management 3 3
Labor Relations 1 1
Trade Relations 2 1
COMPANY TOTAL 17 11
Table 2 : Service
Deliveries on Time 3 3
Handling of Problems 3 3
Technical Assistance 3 2
SERVICE TOTAL 9 8
Table 3 : Products
Quality 4 2
Price 1 2
Packaging 2 1
PRODUCTS TOTAL 7 5
TABLE 4 : SALES
Product knowledge 1 2
Sales Calls 1 1
Sales Service 3 2
SALES TOTAL 5 5

Weighted score of Kay Corp = Weights of each factor * Total rating of factors

Weights for company, service, products, sales are 0.2,0.2,0.4,0.2 respectively.

Weighted score of Kay corp = 0.2*17 + 0.2*9 + 0.4*7 + 0.2*5

= 9

Weighted score of Donna Inc = 0.2*11 + 0.2*8 + 0.4*5 + 0.2*5 = 6.8

So, the weighted score of the buyer, Kay Corp is greater than that of Donna Inc. Hence, it is ideal to choose Kay Corp as the vendor for Eynan enterprises.